Project

General

Profile

20151201 » History » Version 9

Boyd, Steven, 01 December 2015 17:02

1 1 Boyd, Steven
Commissioning and Ops Scheduling meeting : 1/12/2015
2
3
To be held in CR10 in the ATLAS Building @ 15:30 pm 
4
5
Phone access :  ReadyTalk number 
6
US 866-740-1260
7
UK 0800-496-0576
8
Swiss 0800-70-53-51
9
Italy 800182592
10
Access Code: 2155402
11
12
h2. Agenda:
13
14
* Review of tasks and activities identified at last meeting
15 2 Boyd, Steven
** Focus coil progress and soak test
16 1 Boyd, Steven
** Decay Solenoid PSU Repair
17 2 Boyd, Steven
** Hydrogen/Vacuum system work
18
* Upcoming activities : chaired by Victoria/Paul
19 3 Blackmore, Victoria
** "This week's MICE Hall work plan":http://micewww.pp.rl.ac.uk/attachments/download/5295
20 2 Boyd, Steven
** Focus coil operation
21 1 Boyd, Steven
** Hydrogen/Vacuum system installation and commissioning
22 3 Blackmore, Victoria
** Hydrogen alarm system work (requires roof access Tues--Fri)
23 4 Boyd, Steven
** Hall investigation for visual survey of X-ray leakage and reflection paths : Wed 10:30 am - 11:30 am (requires roof access)
24 3 Blackmore, Victoria
** Nitrogen system work (Tues--Thurs)
25
** North-side handrail installation (Wed am)
26 5 Boyd, Steven
* Operations plan 
27 7 Boyd, Steven
** Thursday 3rd (2 pm - 10 pm) : Noon run plan meeting / Close Hall / Beam tune / Prep data-taking / Take TOF 0 calibration (Q789 off) / FC off data / Open Hall
28 9 Boyd, Steven
** Friday 4th ( 12pm - 8 pm ) : 9 am Run plan meeting / Close Hall / Magnetic sweep / Ramp FC to 50 A Flip mode / FC on data-taking / Go to standby on the target overnight
29
** Saturday 5th ( 9 am - 4 pm ) : 9 am Run Plan meeting / 10 am Ramp FC to 100 A / FC-on data taking / Go to standby on the target overnight
30
** Sunday 6th ( 9 am - 4 pm ) : 9 am Run Plan meeting / 10 am Ramp to 150 A / FC-on data-taking / Go to standby on the target overnight
31
** Monday 7th ( 9 am - end ) : 9 am run plan meeting / 10 am Ramp FC to 0 A / Change polarity / Ramp to 100 A / Take data / End and Open Hall
32 5 Boyd, Steven
* Questions
33
** When can we start operations on Friday?
34
** FC expert availability on Sat/Sun (drives when ramps can happen)
35 1 Boyd, Steven
** Hall activities on morning of Monday 7th
36
37
* AOB
38
** [[WorkPlanJanuary2016|To-do list]] for Maintenance shutdown (January 2016)
39 9 Boyd, Steven
40
h2. Minutes
41
42
John C, Ryan, Melissa, Yordan, Celeste, Adam, Paolo, Ed, Durga, Victoria, Vicky, John, Mark, Steve
43
44
Review:
45
46
* Focus coil ramped to 200 A over weekend and then to 225 A and then down. Looked stable.
47
* Decay Solenoid PSU on transit. In Germany but has not arrive at FUG yet.
48
* Hydrogen/Vacuum : Phil is in the Hall
49
50
Upcoming activities:
51
52
* Focus coil operation went well so don't need further commissioning. 
53
  John : did we lose any He on ramp-down.
54
  Mark : about 4-5 %. We're at 68% now.
55
  Vicky : only boiled off during ramp-down 
56
  John : should be OK over running period. 60% is the "safe" level. 
57
* Hydrogen/Vacuum system work on-going. Will take him until Thursday, leaving @ 12 on Thursday
58
* John wants to work on the ceiling inside the Hall.  Hope this will finish on Thursday.
59
* Tim Stanley doing a walk-around on Wednesday mornig
60
* Nitrogen system work : installation nitrogen 
61
* Handrail installed on Wednesday morning.
62
* Two leaks has been found in the Hydrogen system. An external leak which can be fixed "easily". The major leak
63
  is on the inside of the bore in the magnet. Suspicion is that the leak is related to Indium o-rings but this can't
64
  be fixed from outside. We either need to change the medium or take the magnet apart to get at the leak.  Plan to 
65
  fix during January.
66
  Chris: can we do different absorbers now?
67
  Josef : yes. Could run methanol or water now if necessary. If we put in pressure we might collapse the end bore 
68
  a bit. 
69
  Mike : we could put methanol in there. It would have to be run past the safety committee. 
70
  John : how would we get the methanol out.
71
  Mike : let it evaporate once we take the FC out. 
72
  Ed : is there any possible risk in running in a mode we have not studied?
73
  Mike : need to think this through in the safety committee. 
74
  Procedure : Steve W should drive this in consultation with the safety committee.
75
  Mike will get the ball rolling on getting this through the SC with the hope that we can run from the 14th to 18th.
76
  Chris will decide on methanol vs ethanol
77
  John will buy it in.
78
  Timescale to know if we're going to do this is one week. Next meeting is a decision point.
79
  John suggests removing the FC and putting a chunk of plastic in the beam? 
80
  Ken doesn't think we should do methanol. Or water. Is there a gas which wouldn't freeze but would liquify? Argon?
81
  John says that you want a material which will have 5-10% of X0 in absorber?  Pour in LN2 without the cooler? 
82
  
83
  * Chris will look for "best" material and return a short list of possibilities by Thursday night.
84
  * Absorber team to look at late this week & into next week
85
86
87
* Mike wants us to document the polarity change properly
88
89
AOB:
90
91
* Josef: Do we want to press for the ECE coil at the magnet review?
92
  Victoria : would prefer ECE running